OpenStudy (ammarah):

Graph (x+4)^2=-8(y-2) Parabola? Vertex: (-4,2) Directrix:y=2?? Please check my answers @ganeshie8

ganeshie8 (ganeshie8):

its a parabola cuz you have only one square term

ganeshie8 (ganeshie8):

vertex is right

ganeshie8 (ganeshie8):

directrix looks wrong

ganeshie8 (ganeshie8):

check again

ganeshie8 (ganeshie8):

\(\large (x+4)^2=-8(y-2)\) compare wid standard form : \(\large (x-h)^2=4p(y-k)\)

ganeshie8 (ganeshie8):

\(\large 4p = -8 \implies p = -2\) right ?

OpenStudy (ammarah):

yes y=-2?

ganeshie8 (ganeshie8):

Directrix : \(\large y = k-p\)

OpenStudy (ammarah):

-4

ganeshie8 (ganeshie8):

you should get : Directrix : \(\large y = 2 --2 \implies y = 4\)

OpenStudy (ammarah):

ok can u show me how the graph looks like....idk what points to plug in

ganeshie8 (ganeshie8):

just sketch it roughly

ganeshie8 (ganeshie8):

first plot vertex and directrix

OpenStudy (ammarah):



ganeshie8 (ganeshie8):

Excellent ! and you must be knowing that the parabola always opens toward Focus. and Focus will be on mirror opposite side of directrix

OpenStudy (ammarah):

how do u find the focus?

ganeshie8 (ganeshie8):